# Revisiting Optimal Proportions for Binary Responses: Insights from Incorporating the Absent Perspective of Type-I Error Rate Control

### v2

@@ Abstract
This work revisits optimal response-adaptive designs from a type-I error rate perspective, highlighting when and how much these allocations exacerbate type-I error rate inflation — an issue previously undocumented. We explore a range of approaches from the literature that can be applied to reduce type-I error rate inflation. However, we found that all of these approaches fail to give a robust solution to the problem. To address this, we derive two optimal proportions, incorporating the more robust score test (instead of the Wald test) with finite sample estimators (instead of the unknown true values) in the formulation of the optimization problem. One proportion optimizes statistical power and the other minimizes the total number failures in a trial while maintaining a predefined power level. Through simulations based on an early-phase and a confirmatory trial we provide crucial practical insight into how these new optimal proportion designs can offer substantial patient outcomes advantages while controlling type-I error rate. While we focused on binary outcomes, the framework offers valuable insights that naturally extend to other outcome types, multi-armed trials and alternative measures of interest.
Keywords Neyman Allocation; RSHIR Allocation; Wald Test; Type-I Error Rate; Patient Benefit.
@@ 1 Introduction
While adaptive designs hold promise, their practical application in confirmatory trials lags behind their theoretical prominence (Burnett et al., 2020; Dimairo et al., 2015). Response-adaptive randomization (RAR) designs is an example of this phenomena, having a history spanning nearly a century (Thompson, 1933) but remaining more a topic of persistent controversy among statisticians (Robertson et al., 2023) than a practically chosen design option. This is partly due to the need to balance trade-offs in terms of ethical considerations, power, and type-I error rate as part of achieving acceptability for their practical application. Often discussions of the use of RAR have focused on the trade-off between patient-benefit metrics and statistical power (type-II error) while type-I error rate has been widely ignored in the theoretical literature. From a regulatory perspective type-I error rate control is an absolute minimal requirement for any trial design without which concerns about power or patient-benefit become tangential.
In this paper, we provide new insights for clinical trial designers by proposing two novel optimal RAR designs that explicitly consider type-I error rate control alongside power and ethical benefits. Our focus is on two-armed trials with binary endpoints, which are commonly used in clinical research. In a literature review on PubMed, we identified over 2,200 papers that discuss Phase III trials with a binary endpoint published in the last 10. We reexamine from a new perspective the paper by Rosenberger et al. (2001). We reexamine from a new perspective the paper by Rosenberger et al. (2001), who first proposed an approach based on targeting optimal allocation proportions.
In Section 2, we present two optimal allocation proportions that were presented in Rosenberger et al. (2001) and serve as the foundation of our analysis. The first, Neyman allocation, maximizes the statistical power of a widely used Wald-type test. The second, RSHIR allocation, an abbreviation of the names of authors of Rosenberger et al. (2001), minimizes the expected number of patient failures during the trial while ensuring the test maintains a predefined power. Both allocations are derived for the Wald test. Other have reported issues with its use as an inference method following response-adaptive trials (Antognini et al., 2018). Here, we reveal a novel finding: using Neyman or RSHIR allocation further exacerbates type-I error rate inflation to a degree that may surprise readers. To our knowledge, this issue has not been documented in the prior literature.
@@ 2 Notation and Framework for Optimal Response Adaptive Randomization
In this and the next two sections, we consider a two-armed binary trial with a fixed sample size $n=n_{0}+n_{1}$, where $n_{0}$ and $n_{1}$ are the sample sizes for the control and treatment arms, respectively. The responses of individual patients are modeled as random variables $Y_{ki}$. Here, $k\in\{0,1\}$ indicates the treatment ($k=1$) or control ($k=0$), and $i\in\{1,\dots,n\}$ indexes the patients. For each arm $k\in\{0,1\}$, the random variables $Y_{k1},\dots,Y_{kn_{k}}$ are assumed to be independent and identically distributed (i.i.d.) as Bernoulli random variables
$Y_{ki}\mathop{\sim}\limits^{\mathrm{iid}}\text{Bern}(p_{k}),\quad i=1,\dots,n_{k};\quad k=0,1;\quad n_{0}+n_{1}=n,$ (1)
@@
Out of the $2\times n$ potential outcomes, only $n$ are observed in the trial. The treatment assignment for patient $i$ is indicated by a binary variable $A_{i}$, where $A_{i}=1$ if the patient is assigned to the experimental arm and $A_{i}=0$ otherwise. The probability of assigning a patient to treatment $k$ is therefore $P(A_{i}=1)$, and $P(A_{i}=0)=1-P(A_{i}=1)$. Because randomization minimizes bias, reduces confounding, and ensures comparability between treatment groups, we focus here only on designs that satisfy: $P(A_{i}=k)>0$ for all $i=1,\dots,n$ and $k=0,1$. The simplest (non-adaptive) randomization scheme is complete randomization (CR), where $P(A_{i}=k)=1/2$ for all $i$ and $k$ in a two-arm trial. CR attains an equal allocation proportion (on average) and it is often assumed to maximize statistical power (Berger et al., 2021; Friedman et al., 2015) and control the type-I error rate.
@@
At the trial’s conclusion, a statistical test is chosen for final inference. The two-sample Wald test is widely used in RAR literature for evaluating optimal designs (Hu and Rosenberger, 2006). This test considers a measure of interest based on the success probabilities, denoted as $f(p_{0},p_{1})$. In this paper, we focus (as does Rosenberger et al. (2001)) on the mean parameter difference $f(p_{0},p_{1})=p_{1}-p_{0}$, though other measures such as the log relative risk or log odds ratio can also be used as suggested by Rosenberger et al. (2001). . The general null hypothesis is
@@
$Z_{1}=\frac{\hat{p}_{1}-\hat{p}_{0}}{\sqrt{\frac{\hat{p}_{0}\hat{q}_{0}}{n_{0}}+\frac{\hat{p}_{1}\hat{q}_{1}}{n_{1}}}}.$ (3)
@@ 2.1 Two Optimal Proportions for the Wald Test
The first allocation we are discussing is Neyman allocation based on Neyman (1934); Robbins (1952). The aim is to maximize the statistical power of the Wald test $Z_{1}$, i.e. the probability to reject the null hypothesis when $p_{0}\neq p_{1}$. Maximizing power of the test is equivalent to minimizing the variance of the test statistic
$s_{T}^{2}=\frac{p_{0}q_{0}}{n_{0}}+\frac{p_{1}q_{1}}{n_{1}},$ (4)
Hence, the optimization problem can be formulated as
@@
$\rho_{N_{1}}=\frac{\sqrt{p_{1}q_{1}}}{\sqrt{p_{0}q_{0}}+\sqrt{p_{1}q_{1}}},$ (5)
@@
$\rho_{R_{1}}=\frac{\sqrt{p_{1}}}{\sqrt{p_{0}}+\sqrt{p_{1}}}.$ (6)
@@ 2.2 Targeting Procedures
Two more recently developed methods are the doubly adaptive biased coin design (DBCD), introduced by Eisele (1994) and Eisele and Woodroofe (1995) and the efficient randomized-adaptive design (ERADE) introduced by Hu et al. (2009). ERADE is the discretized adaptation of Hu and Zhang’s DBCD aiming to decrease the variance of the design. Given a parameter $\alpha\in(0,1)$, the probability of assigning patient $j+1$ to treatment 1 is determined as follows:
$p_{1}(n_{1}(j),\hat{p}_{1}(j))=\begin{cases}\alpha\hat{p}_{1}(j),&\text{if $n_{1}(j)/j>\hat{p}_{1}(j)$,}\\ \hat{p}_{1}(j),&\text{if $n_{1}(j)/j=\hat{p}_{1}(j)$,}\\ 1-\alpha(1-\hat{p}_{1}(j)),&\text{if $n_{1}(j)/j<\hat{p}_{1}(j)$.}\end{cases}$ (7)
Following the recommendation of Burman (1996), the parameter $\alpha$ is typically chosen within the range of $0.4$ to $0.7$. We choose $\alpha=0.5$.
@@ 2.3 Simulation Results
In this section, we present simulation results for the Wald test under three allocation rules: CR ($\rho_{CR}$), Neyman ($\rho_{N_{1}}$), and RSHIR ($\rho_{R_{1}}$), using ERADE as the targeting procedure. A significance level of $\alpha=0.05$ was used for the Wald test and will be used for all future tables and figures in this paper. The simulations in this section were conducted with a minimal burn-in period of $B=4$, two patients per arm allocated using permuted block randomization, and each scenario was simulated 10,000 times, ensuring a Monte Carlo error of less than 0.5% for both power and type-I error rate (Morris et al., 2019). Table 1 presents the results for type-I error rate across the entire parametric space. Additionally, it reports results for a fixed $p_{0}=0.2$ ($p_{0}=0.7$) under an increasing (decreasing) sequence of treatment effects to illustrate differences in power.
@@
Table 1: Power or Type-I Error Rate, proportion allocated to the treatment arm $n_{1}/n$, and expected number of successes (ENS) for $n=50$ and different settings of $p_{0}$ and $p_{1}$.
@@
Figure 1: Comparison of Type-I error rates of: the Neyman proportion testing with the Wald test $\rho_{N_{1}}$, the RSHIR proportion testing with the Wald test $\rho_{N_{1}}$, complete randomization testing with the Wald test $\rho_{CR_{1}}$, complete randomization testing with the score test $\rho_{CR_{0}}$, the Neyman-like proportion testing with the score test $\rho^{n}_{N_{0}}$ and the RSHIR-like proportion testing with the score test $\rho^{n}_{R_{0}}$.
@@ 3.1 Agresti & Caffo Correction
Table 2: Comparison of type-I error rate or power and ENS for various settings of $p_{0}$ and $p_{1}$: The tables provide results for the Agresti & Caffo correction and differnt burn-in periods ($B=12$, $B=30$, $B=120$) for two sample sizes, $n=50$ (upper table) and $n=200$ (lower table).
@@ 3.2 Ensuring Existence of Variance Estimators
Table 3: Results for sampling with equal probability if one of the two variance estimators is 0. Power or Type-I Error Rate, proportion allocated to the treatment arm $n_{1}/n$, and expected number of successes (ENS) for for $n=50$ and different settings of $p_{0}$ and $p_{1}$.
@@ 3.3 Burn-In Periods to Stabilize Estimators and Decrease Adaptivity
A burn-in period is defined as the initial allocation of the first $B\in\{1,\dots,n\}$ patients with equal probability to each arm. To ensure an equal number of patients ($B/2$) in each arm, we enforce balance using methods such as simple random allocation or permuted block randomization. This step stabilizes the MLEs before adapting the allocation probabilities to reduce the likelihood of assigning patients in the wrong direction. The minimal burn-in size for RSHIR is one patient per arm ($B=2$), whereas for Neyman, it is two patients per arm ($B=4$), see Table 1. Table 2 demonstrates that (as the other approaches) the procedures with using larger burn-in periods still result in inflated type-I error rates. Moreover, increasing the burn-in size comes at the cost of diminishing patient benefit gains (RSHIR) or power gains (Neyman). For Table 2 we used a burn-in period of six patients per arm ($B=12$), as recommend by Rosenberger et al. (2001), and 60% of the total trial size to demonstrate the the asymptotic behavior. In smaller trials, this trade-off often forces a near complete convergence to a CR design in an attempt to control the type-I error rate, though full control may not always be achievable, see Section 3.4. In larger trials, the same burn-in size allows for adaptive sampling while retaining error control, see Table 2.
@@ 3.4 Using the Score Test: Testing with a Pooled Variance Estimator
Eberhardt and Fligner (1977) and Fleiss et al. (1981, page 30) observed that the Wald-type test statistic $Z_{1}$ often exhibits an inflated type-I error rate in non-adaptive settings. To mitigate this issue, they propose testing the null hypothesis $H_{0}:p_{0}=p_{1}$ using the score test statistic defined as
$Z_{0}=\frac{\hat{p}_{1}-\hat{p}_{0}}{\sqrt{\hat{p}\hat{q}\left(\frac{1}{n_{0}}+\frac{1}{n_{1}}\right)}},$ (8)
where $\hat{p}$ is the overall proportion of successes in the trial, calculated as the total number of successes divided by the trial size $n$, and $\hat{q}=1-\hat{p}$. The score test $Z_{0}$, is commonly used (Agresti, 2013) and it can be shown that the squared statistic, $Z_{0}^{2}$, is equivalent to the widely used Pearson chi-squared test.
In contrast, the variance of the test statistic $Z_{1}$ that was used by Tschuprow (1923); Robbins (1952); Rosenberger et al. (2001) incorporates the individual estimates of $p_{0}$ and $p_{1}$ in the estimator of the variance of the test statistic, enabling the construction of confidence intervals that align with the test decision. While using the Wald test typically achieves higher power compared to the score test, it does so at the expense of inflating the type-I error rate, even in non-adaptive settings (Fleiss et al., 1981, page 30).
@@
0.2 0.3 12.6% 7.4% 2.5% 0.5 (0) 0.56 (0.1488) 0.58 (0.1391) 12.6 12.8 12.9
0.2 0.5 63.9% 35.5% 22.6% 0.5 (0) 0.62 (0.1337) 0.71 (0.0913) 17.5 19.4 20.8
0.2 0.7 97.2% 73.4% 54.7% 0.5 (0) 0.55 (0.1494) 0.79 (0.0472) 22.5 23.8 29.8
Table 4: Testing with the score test: Power or Type-I Error Rate, proportion allocated to the treatment arm $n_{1}/n$, and expected number of successes (ENS) for for $n=50$ and different settings of $p_{0}$ and $p_{1}$.

### v3

# Revisiting Optimal Allocations for Binary Responses: Insights from Considering Type-I Error Rate Control

@@ Abstract
This work revisits optimal response-adaptive designs from a type-I error rate perspective, highlighting when and how much these allocations exacerbate type-I error rate inflation — an issue previously undocumented. We explore a range of approaches from the literature that can be applied to reduce type-I error rate inflation. However, we found that all of these approaches fail to give a robust solution to the problem. To address this, we derive two optimal allocation proportions, incorporating the more robust score test (instead of the Wald test) with finite sample estimators (instead of the unknown true values) in the formulation of the optimization problem. One proportion optimizes statistical power and the other minimizes the total number failures in a trial while maintaining a fixed variance level. Through simulations based on an early-phase and a confirmatory trial we provide crucial practical insight into how these new optimal proportion designs can offer substantial patient outcomes advantages while controlling type-I error rate. While we focused on binary outcomes, the framework offers valuable insights that naturally extend to other outcome types, multi-armed trials and alternative measures of interest.
Keywords Neyman allocation; patient benefit; RSHIR allocation; score test; Wald test.
@@ 1 Introduction
While adaptive designs hold promise, their practical application in confirmatory trials lags behind their theoretical prominence (e.g., Burnett et al., 2020; Dimairo et al., 2015). Response-adaptive randomization (RAR) designs is an example of this phenomenon, having a history spanning nearly a century (Thompson, 1933) but remaining more a topic of persistent controversy among statisticians (Robertson et al., 2023) than a practically chosen design option. This is partly due to the need to balance trade-offs in terms of patient-benefit considerations, power, and type-I error rate as part of achieving acceptability for their practical application. Often discussions of the use of RAR have focused on the trade-off between patient-benefit metrics and statistical power while the type-I error rate has been ignored in the theoretical literature. From a regulatory perspective, type-I error rate control is an absolute minimal requirement for any trial design, without which concerns about power or patient benefit become tangential.
In this paper, we provide new insights for clinical trial designers by proposing two novel optimal RAR designs that explicitly consider type-I error rate control alongside power and patient benefits. Our focus is two-armed trials with binary endpoints, which are commonly used in clinical research. In a literature review on PubMed, we identified over 2,200 papers that discuss Phase III trials with a binary endpoint published in the last 10 years. We reexamine from a new perspective the paper by Rosenberger et al. (2001), who first proposed an approach based on targeting optimal allocation proportions, and is one of the most cited papers (335) in the RAR literature.
In Section 2, we present two optimal allocation proportions that were presented in Rosenberger et al. (2001) and serve as the foundation of our analysis. The first, Neyman allocation, maximizes the statistical power of a widely used Wald test. The second, RSHIR allocation, an abbreviation of the names of authors of Rosenberger et al. (2001), minimizes the expected number of patient failures during the trial while ensuring the test maintains a predefined variance level. Both allocations are derived for the Wald test. Other have reported issues with its use as an inferential method following response-adaptive trials (Baldi Antognini et al., 2018). Here, we reveal a novel finding: using Neyman or RSHIR allocation further exacerbates type-I error rate inflation to a degree that may surprise readers. To our knowledge, this issue has not been documented in the prior literature.
@@
Our main contribution is the development of a procedure and test statistic that preserves the type-I error rate for binary response RAR designs. Note that the structure of this paper follows the way in which we arrived at this procedure. We emphasize, that in practice, one should first ensure that type-I error rates are controlled before power and patient-benefit gains are considered.
@@ 2 Notation and Framework for Optimal Response Adaptive Randomization
In this and the next two sections, we consider a two-armed binary trial with a fixed sample size $n=n_{0}+n_{1}$, where $n_{0}$ and $n_{1}$ are the sample sizes for the control and treatment arms, respectively. The responses of individual patients are modeled as random variables $Y_{ki}$. Here, $k\in\{0,1\}$ indicates the treatment ($k=1$) or control ($k=0$), and $i\in\{1,\dots,n\}$ indexes the patients. For each arm $k\in\{0,1\}$, the random variables $Y_{k1},...,Y_{kN}$ are assumed to be independent and identically distributed (i.i.d.) as Bernoulli random variables
$Y_{ki}\mathop{\sim}\limits^{\mathrm{i.i.d.}}\text{Bern}(p_{k}),\quad i=1,\dots,n_{k};\quad k=0,1;\quad n_{0}+n_{1}=n,$
@@
Out of the $2\times n$ potential outcomes (corresponding to two treatments and $n$ patients), only $n$ are observed in the trial, since for each patient we can observe the outcome under only one treatment. The treatment assignment for patient $i$ is indicated by a binary variable $A_{i}$, where $A_{i}=1$ if the patient is assigned to the experimental arm and $A_{i}=0$ otherwise. The probability of assigning a patient to treatment $k$ is therefore $P(A_{i}=1)$, and $P(A_{i}=0)=1-P(A_{i}=1)$. Because randomization minimizes bias, reduces confounding, and ensures comparability between treatment groups, we focus here only on designs that satisfy: $P(A_{i}=k)>0$ for all $i=1,\dots,n$ and $k=0,1$. The simplest (non-adaptive) randomization scheme is complete randomization (CR), where $P(A_{i}=k)=1/2$ for all $i$ and $k$ in a two-arm trial. CR attains an equal allocation proportion (on average) and it is often assumed to maximize statistical power (Berger et al., 2021; Friedman et al., 2015) and control the type-I error rate.
@@
At the trial’s conclusion, a statistical test is chosen for final inference. The two-sample Wald test is widely used in RAR literature for evaluating optimal designs (Hu and Rosenberger, 2006). This test considers a measure of interest based on the success probabilities, denoted as $f(p_{0},p_{1})$. In this paper, we focus (as does Rosenberger et al. (2001)) on the mean parameter difference $f(p_{0},p_{1})=p_{1}-p_{0}$, though other measures such as the log relative risk or log odds ratio can also be used as suggested by Rosenberger et al. (2001). A complete table of optimal allocation proportions is given by Pin et al. (2024, p. 325). The general null hypothesis is
@@
$Z_{1}=\frac{\hat{p}_{1}-\hat{p}_{0}}{\sqrt{\frac{\hat{p}_{0}\hat{q}_{0}}{n_{0}}+\frac{\hat{p}_{1}\hat{q}_{1}}{n_{1}}}}.$
@@ 2.1 Two Optimal Proportions for the Wald Test
The first allocation we discuss is Neyman allocation (Neyman, 1934; Robbins, 1952). The aim is to fix the asymptotic variance of the the treatment difference estimator (or equivalently the inverse of the Fisher information, as in Jennison and Turnbull’s formulation (2000, p. 327)), given by $s_{T}^{2}=(p_{0}q_{0})/n_{0}+(p_{1}q_{1})/n_{1}$ (Rosenberger et al., 2001, pp. 910), and under that constraint optimize
@@
$\rho_{N_{1}}=\frac{\sqrt{p_{1}q_{1}}}{\sqrt{p_{0}q_{0}}+\sqrt{p_{1}q_{1}}},$
@@
$\rho_{R_{1}}=\frac{\sqrt{p_{1}}}{\sqrt{p_{0}}+\sqrt{p_{1}}}.$
@@ 2.2 Targeting Procedures
Two more recently developed methods are the doubly adaptive biased coin design (DBCD), introduced by Eisele (1994); Eisele and Woodroofe (1995) and modified by Hu and Zhang (2004); and the efficient randomized-adaptive design (ERADE) introduced by Hu et al. (2009). ERADE is the discretized adaptation of the doubly adaptive biased coin design (DBCD) by Eisele (1994) aiming to decrease the variance of the design. Given a parameter $\alpha\in(0,1)$, the probability of assigning patient $j+1$ to treatment 1 is determined as follows:
$\rho(n_{1}(j),\hat{\rho}(j))=\begin{cases}\alpha\hat{\rho}(j),&\text{if $n_{1}(j)/j>\hat{\rho}(j)$,}\\ \hat{\rho}(j),&\text{if $n_{1}(j)/j=\hat{\rho}(j)$,}\\ 1-\alpha(1-\hat{\rho}(j)),&\text{if $n_{1}(j)/j<\hat{\rho}(j)$,}\end{cases}$
where $\hat{\rho}(j)$ is an estimated optimal proportion and $n_{1}(j)$ the number of patients allocated to arm 1 after the allocation of $j$ patients. Following the recommendation of Burman (1996), the parameter $\alpha$ is typically chosen within the range of $0.4$ to $0.7$. We choose $\alpha=0.5$.
@@ 2.3 Simulation Results
In this section, we present simulation results for the Wald test under three allocation rules: CR ($\rho_{CR}$), Neyman ($\rho_{N_{1}}$), and RSHIR ($\rho_{R_{1}}$), using ERADE as the targeting procedure. A significance level of $\alpha=0.05$ was used for the Wald test and will be used for all future tables and figures in this paper. The simulations in this section were conducted with a minimal burn-in period of $B=4$, two patients per arm allocated using permuted block randomization, and each scenario was simulated 10,000 times, ensuring a Monte Carlo error of less than 0.5% for both power and type-I error rate (Morris et al., 2019). We use a burn-in of 2 patients per arm consistently throughout the paper. There is no consensus in the literature on the optimal burn-in size for response-adaptive trials (Tang et al., 2025). We intentionally use the minimal burn-in (2 patients per arm) to investigate the maximal extent of type-I error rate inflation.
Table 1 presents the results for type-I error rate across the entire parametric space. Additionally, it reports results for a fixed $p_{0}=0.2$ ($p_{0}=0.7$) under an increasing (decreasing) sequence of treatment effects to illustrate differences in power.
@@
Table 1: Power or Type-I Error Rate, proportion allocated to the treatment arm $n_{1}/n$, and expected number of successes (ENS) for $n=50$, a burn-in size of $2$ patients per arm and different settings of $p_{0}$ and $p_{1}$.
@@
Figure 1: Comparison of Type-I error rates of: the Neyman proportion testing with the Wald test $\rho_{N_{1}}$, the RSHIR proportion testing with the Wald test $\rho_{N_{1}}$, complete randomization testing with the Wald test $\rho_{CR_{1}}$, complete randomization testing with the score test $\rho_{CR_{0}}$, the Neyman-like proportion testing with the score test $\rho^{n}_{N_{0}}$ and the RSHIR-like proportion testing with the score test $\rho^{n}_{R_{0}}$. In the both figures a burn-in size of $2$ patients per arm was used.
@@ 3.1 Agresti & Caffo Correction
Table 2: Comparison of type-I error rate or power and ENS for various settings of $p_{0}$ and $p_{1}$: The tables provide results for the Agresti & Caffo correction and different burn-in periods ($B=12$, $B=30$, $B=120$) for two sample sizes, $n=50$ (upper table) and $n=200$ (lower table). For the Agresti & Caffo correction a minimal burn of size $2$ patients per arm was used.
@@ 3.2 Ensuring Existence of Variance Estimators
Table 3: Results for sampling with equal probability if one of the two variance estimators is 0. Power or Type-I Error Rate, proportion allocated to the treatment arm $n_{1}/n$, and expected number of successes (ENS) for $n=50$, a burn-in size of $2$ patients per arm and different settings of $p_{0}$ and $p_{1}$.
@@ 3.3 Burn-In Periods to Stabilize Estimators and Decrease Adaptivity
A burn-in period is defined as the initial allocation of the first $B\in\{1,\dots,n\}$ patients with equal probability to each arm. To ensure an equal number of patients ($B/2$) in each arm, we enforce balance using methods such as simple random allocation or permuted block randomization. This step stabilizes the MLEs before adapting the allocation probabilities to reduce the likelihood of assigning patients in the wrong direction. The minimal burn-in size for RSHIR is one patient per arm ($B=2$), whereas for Neyman, it is two patients per arm ($B=4$), see Table 1. Table 2 demonstrates that (as the other approaches) the procedures with using larger burn-in periods still result in inflated type-I error rates. Moreover, increasing the burn-in size comes at the cost of diminishing patient-benefit gains (RSHIR) or power gains (Neyman). For Table 2 we used a burn-in period of six patients per arm ($B=12$), as recommend by Rosenberger et al. (2001), and 60% of the total trial size to demonstrate the the asymptotic behavior. In smaller trials, this trade-off often forces a near complete convergence to a CR design in an attempt to control the type-I error rate, though full control may not always be achievable, see Section 3.4. In larger trials, the same burn-in size allows for adaptive sampling while retaining error control, see Table 2.
@@ 3.4 Using the Score Test: Testing with a Pooled Variance Estimator
Eberhardt and Fligner (1977) and Fleiss et al. (1981, page 30) observed that the Wald-type test statistic $Z_{1}$ often exhibits an inflated type-I error rate in non-adaptive settings. To mitigate this issue, they propose testing the null hypothesis $H_{0}:p_{0}=p_{1}$ using the score test statistic (Agresti, 2002, p. 14), which is defined as
$Z_{0}=\frac{\hat{p}_{1}-\hat{p}_{0}}{\sqrt{\hat{p}\hat{q}\left(\frac{1}{n_{0}}+\frac{1}{n_{1}}\right)}},$
where $\hat{p}$ is the overall proportion of successes in the trial, calculated as the total number of successes divided by the trial size $n$, and $\hat{q}=1-\hat{p}$. Comparing the CR results from Table 1 with those from Table 4 and Figure 1 confirms that this test provides superior control over the type-I error rate.
The score test $Z_{0}$, is commonly used (Agresti, 2002) and it can be shown that the squared statistic, $Z_{0}^{2}$, is equivalent to the widely used Pearson chi-squared test (Scott, 2020, p. 119). In contrast, the variance of the test statistic $Z_{1}$ that was used by Tschuprow (1923); Robbins (1952); Rosenberger et al. (2001) incorporates the individual estimates of $p_{0}$ and $p_{1}$ in the estimator of the variance of the test statistic, enabling the construction of confidence intervals that align with the test decision. While using the Wald test typically achieves higher power compared to the score test, it does so at the expense of inflating the type-I error rate, even in non-adaptive settings (Fleiss et al., 1981, page 30).
@@
0.2 0.3 12.6% 7.4% 2.5% 12.2% 10.4% 0.5 (0) 0.56 (0.1488) 0.58 (0.1391) 0.46 (0.0023) 0.47 (0.0011) 12.6 12.8 12.9 12.3 12.4
0.2 0.5 63.9% 35.5% 22.6% 60.2% 61.1% 0.5 (0) 0.62 (0.1337) 0.71 (0.0913) 0.43 (0.0016) 0.50 (0.0020) 17.5 19.4 20.8 16.5 17.6
0.2 0.7 97.2% 73.4% 54.7% 96.6% 96.1% 0.5 (0) 0.55 (0.1494) 0.79 (0.0472) 0.46 (0.0023) 0.60 (0.0050) 22.5 23.8 29.8 21.5 25.1
Table 4: Comparison of five testing procedures using the score test: Type-I Error Rate or Power, proportion allocated to the treatment arm $n_{1}/n$, and expected number of successes (ENS) for $n=50$ and a burn-in size of $2$ per arm.